imqwerty (imqwerty):

prove that \[3^{2^{n}}-1\] can be written as the sum of two squares for all positive integers n.

OpenStudy (kainui):

\((3^{2^{n-1}})^2\) and \(i^2\) are both perfect squares, solved! :^)

imqwerty (imqwerty):

lol yes correct :D

imqwerty (imqwerty):

today ima post lots of ques like this (:

imqwerty (imqwerty):

better ones xD

OpenStudy (kainui):

Ok but seriously this is as far as I've gotten: \[(3^{2^{n-1}}+1)(3^{2^{n-1}}-1)=a^2+b^2\] I feel like I've seen this sorta thing before, gotta think how to continue (if this is even a step in the right direction).

ganeshie8 (ganeshie8):

It is sufficient to show that none of the prime divisors of \(3^{2^n}-1\) are of form \(4k+3\)

ganeshie8 (ganeshie8):

but since this is supposed to be trick questions, im not gonna attempt... :)

OpenStudy (kainui):

Ok time for me to make a serious second attempt at internalizing Fermat's little theorem I guess.

OpenStudy (kainui):

Well maybe I'm just assuming that's going to be useful here, beats me.

ganeshie8 (ganeshie8):

the prime factorization has a really interesting pattern http://www.wolframalpha.com/input/?i=Table%5Bfactor+3%5E2%5En-1%2C%7Bn%2C1%2C6%7D%5D for the first 6 terms that i wolfram gave, each subsequent term is giving a new prime number

ganeshie8 (ganeshie8):

2^3, 2^4×5, 2^5×5×41, 2^6×5×17×41×193, 2^7×5×17×41×193×21523361, 2^8×5×17×41×193×21523361×926510094425921

ganeshie8 (ganeshie8):

except for 2, all the prime factors are of form 4k+1 so we can represent the terms as sum of two squares..

OpenStudy (kainui):

Why is this true? \(\color{blue}{\text{Originally Posted by}}\) @ganeshie8 It is sufficient to show that none of the prime divisors of \(3^{2^n}-1\) are of form \(4k+3\) \(\color{blue}{\text{End of Quote}}\)

ganeshie8 (ganeshie8):

that is due to fermat https://www.math.hmc.edu/funfacts/ffiles/20008.5.shtml

ganeshie8 (ganeshie8):

if any prime factors are of form 4k+3, then they must come in even powers... so we're good if we somehow show that none of the prime divisors are of form 4k+3

ganeshie8 (ganeshie8):

the pattern just keeps going with a new prime of form \(4k+1\) with each subsequent term : \[a_n = 2p_na_{n-1}\]

ganeshie8 (ganeshie8):

conjecture : every odd prime divisor of \(3^{2^n}-1\) is of form \(4k+1\)

OpenStudy (kainui):

Interesting, so there could still be prime factors of the form 4k+3 as long as they occur in pairs, makes sense but not sure how to show this. Gonna grab some more coffee and think about this.

OpenStudy (kainui):

Wait, how sure are you of this pattern, or is this just a guess? \(\color{blue}{\text{Originally Posted by}}\) @ganeshie8 the pattern just keeps going with a new prime of form \(4k+1\) with each subsequent term : \[a_n = 2p_na_{n-1}\] \(\color{blue}{\text{End of Quote}}\)

imqwerty (imqwerty):

yes thats correct :) this can help understand - http://prntscr.com/928baa

ganeshie8 (ganeshie8):

yes that is just a guess @Kainui ... i don't have any proofs yet..

ganeshie8 (ganeshie8):

i have checked the first 11 terms and saw that pattern... but i have a strong feeling that the pattern must break at some point

OpenStudy (kainui):

Hahaha ok ok, I am just in unfamiliar territory so I wasn't sure where the line is between stuff you know and stuff you're guessing at. I am noticing though we have this sorta recursion pattern going, \[(3^{2^{n+1}}-1) = (3^{2^{n}}-1)*a_n \] I don't know if there's anything we can say about \(a_n=3^{2^n}+1\), but basically our number is a product of these \(a_n\) values.

ganeshie8 (ganeshie8):

yeah that also means \(a_i\mid a_j\) for all \(i\le j\)

OpenStudy (kainui):

Cute form: \[3^{2^n}-1 = 2 \prod_{k=1}^n (3^{2^k}+1)\]

OpenStudy (anonymous):

aww

ganeshie8 (ganeshie8):

Cute form: \[3^{2^n}-1 = 2 \prod_{k=0}^{\color{red}{n-1}} (3^{2^k}+1)\]

OpenStudy (kainui):

Redemption alternate cute form: \[3^{2^n}-1 = 2 \sum_{k=0}^{2^n-1} 3^k\]

ganeshie8 (ganeshie8):

that looks nice but sum may not be so useful when analysing prime factor though...

ganeshie8 (ganeshie8):

Redemption alternate cute form: \[3^{2^n}-1 = 2 \sum_{k=0}^{2^{\color{red}{n-1}}} 3^k\] (just nitpicking :) )

OpenStudy (kainui):

Wait that's not right though is it?

ganeshie8 (ganeshie8):

you're just expanding using \(x^n-y^n = (x-y)(stuff)\) right ?

OpenStudy (kainui):

No, this is the geometric series

ganeshie8 (ganeshie8):

they are same..

ganeshie8 (ganeshie8):

\[x^n-y^n = (x-y)(x^{n-1}+x^{n-2}y + \cdots + y^{n-1})\]

ganeshie8 (ganeshie8):

that identity follows from geometric series yeah :)

OpenStudy (kainui):

\[\frac{x^a - 1}{x-1} = \sum_{k=0}^{a-1} x^k\] \(x=3\), \(a=2^n\)

ganeshie8 (ganeshie8):

it is a mystery why that sequence is laying one prime with each subsequent term this cannot be true, it should break at some point...

OpenStudy (kainui):

\[3^{2^n}-1 = 2 \sum_{k=0}^{2^n-1} 3^k\] If we think back to how all Mersenne primes were of the form of \(2^p-1\) because if p isn't prime, for numbers like p=6 we would have in binary all the factors of 6 have representations: 1*111111=11*10101=111*1001 Similarly I think we can look at this representation in base 3 and see that all its prime factors come in pairs because factors come in pairs.

OpenStudy (kainui):

So what that means is that at least saves it because them coming in pairs means \(3^2 \equiv 1 \mod 4\) is maintained... Which doesn't really prove the conjecture either way haha.

OpenStudy (kainui):

So maybe it turns out for higher numbers there are some 4n+3 primes and it doesn't just multiply in a single new prime every time... I don't know.

ganeshie8 (ganeshie8):

is it easy to prove that \(p\mid a_n \implies p\mid a_{n+1}\) ? does that follow from your earlier representation of \(3^{2^n}-1\) as a product ?

OpenStudy (kainui):

Actually there's a funny thing going on here when there are a prime number of powers of 3 in that sum, that means \(2^n-1\) is prime... which means n is prime... lol I'm not really sure I'm not actually going anywhere anymore I'm just sorta exploring around.

ganeshie8 (ganeshie8):

Nvm, it trivially follows from earlier statement \(a_{n+1} = 3^{2^{n+1}}-1 = (3^{2^n}+1)*a_{n}\) if \(p\mid a_{n}\), we do get \(p\mid a_{n+1}\)

OpenStudy (kainui):

Yes what you have just written is true, but I think the way I did it was the + terms were the \(a_n\), so just saying that to make sure you're not mixing anything up, I think this got sorta confusing. Well at least I'm confused lol

ganeshie8 (ganeshie8):

since our goal is to prove that every prime factor, \(p\), is of form \(4k+1\), I guess below might work : If we "somehow" show that \(4\) is the order of \(3^{2^{n-2}}\) in modulus \(p\), then we can argue that the order must divide \(\phi(p)\). \(4\mid \phi(p) \) \(4\mid p-1\) \(p=4k+1\)

ganeshie8 (ganeshie8):

Got it ! new prime(s) of \(a_n\), that are not occured in the previous terms will be of form \(2^na+1\) (il post a proof for above statement shortly ) clearly \( 2^na+1\equiv 1\pmod{4}\) so we're done !

OpenStudy (kainui):

Ohhhh nice!

ganeshie8 (ganeshie8):

(proof) every new prime divisor of \(a_n\) is of form \(2^na+1\) \(a_{n} = 3^{2^n}-1 = (3^{2^{n-1}}+1) (3^{2^{n-1}}-1)= (3^{2^{n-1}}+1)*a_{n-1}\) New prime(s) occur in the expression \(3^{2^{n-1}}+1\). So just consider this. Let \(p\) be any "odd" prime divisor of \(3^{2^{n-1}}+1\). \(3^{2^{n-1}}+1\equiv 0 \pmod{p}\) \(3^{2^{n-1}}\equiv -1 \pmod{p}\) \(3^{2^n}\equiv 1 \pmod{p}\) that means \(2^n\)is the order of \(3\) in modulus \(p\). it follows \(2^n \mid \phi(p)\) \(2^n\mid p-1\) \(p=2^na+1\) for some integer \(a\)

ganeshie8 (ganeshie8):

Overall we have proved that the odd prime divisors of \(3^{2^n}-1\) are of form \(4k+1\)

OpenStudy (kainui):

Awesome, this is giving me some more ideas. Also, I am trying to see a little more about these prime divisors and what they are, it looks like we have: \[10 \equiv 0 \mod p\] Is this at all surprising to us?

OpenStudy (kainui):

Awesome, this is giving me some more ideas. Also, I am trying to see a little more about these prime divisors and what they are, it looks like we have: \[10 \equiv 0 \mod p\] Is this at all surprising to us?

ganeshie8 (ganeshie8):

\(10\equiv 0\) ?

OpenStudy (kainui):

Let \(p\) be any "odd" prime divisor of \(3^{2^{n-1}}+1\). \(3^{2^{n-1}}\equiv -1 \pmod{p}\) \(3^{2^n}\equiv 1 \pmod{p}\) Sorry I forgot to put this, I'm pretty tired: \(1\equiv 3^{2^n} \equiv 9*3^{2^{n-1}}\equiv 9*(-1) \pmod{p}\)

ganeshie8 (ganeshie8):

that is true only in one direction, they are not independent congruences...

ganeshie8 (ganeshie8):

\(3^{2^{n-1}}\equiv -1 \pmod{p} \implies 3^{2^n}\equiv 1 \pmod{p}\) not the other way...

OpenStudy (kainui):

Lol ok not sure what that means I'll sleep on it I guess

ganeshie8 (ganeshie8):

we can say \(a \equiv b \pmod{n} \implies a^2\equiv b^2\pmod {7}\) but the converse isn't true we cannot take the statement \(a^2\equiv b^2\pmod{n}\) and plugin \(a=b\)

ganeshie8 (ganeshie8):

\(n\mid (a^2-b^2)\) \(n\mid (a+b)(a-b)\) \(n\) need not be a divisor of \(a-b\)

ganeshie8 (ganeshie8):

Also \(3^{2^n}\) and \(9*3^{2^{n-1}}\) are not same things

imqwerty (imqwerty):

shuld i give some hint?

imqwerty (imqwerty):

\[Solution-->\] i just factored it\[3^{2^n}-1=(3^{2^{n-1}}+1^2).......(3^8+1^2)(3^4+1^2)(3^2+1^2)(3^2-1^2)\]\[3^{2^n}-1=(3^{2^{n-1}}+1^2).......(3^8+1^2)(3^4+1^2)(3^2+1^2)(2^2+2^2)\] note that the product of sum of 2squares can be written as a sum of 2squares-\[(a^2+b^2)(c^2+d^2)=(ad-bc)^2 +(ac+bd)^2\]hence proved